OpenStudy (anonymous):

the perimeter of a rectangle is 14 feet the length is 1 inch longer than the width.Write a system of linear equations and solve the resulting system. let x be the length and y be the width. 2x+2y=? x=y+? what is the length? what is the width?

OpenStudy (jusaquikie):

do it all in inches to make it easier so 14 ft is 168 inches. 2x+2y = 168 x=y+1 you are going to use both these equations and substitute between them. we want to get down to only 1 variable so we are going to substitute one equation for the other like below. 2(y+1)+2y=168 Now distribute your two and then get all your y's on one side. 2y+2+2y=168 4y+2=168 4y=168-2 4y=166 y=166/4 y=41.5 Now take this back to the original problem and plug it in 2x+2(41.5)=168 now solve for x (which should be 1 more than 41.5) 2x+83=168 2x=85 x=85/2 x=42.5 but we could have just added 1 to y since we had the equation (x=y+1)